OpenStudy (devan):

so how would i do this equation: y+6=-1/6(x+3) ?

OpenStudy (amistre64):

do?

OpenStudy (anonymous):

distribute the -1/6 to get y+6 = -x/6 + -1/2, then subtract 6 from both sides.

OpenStudy (anonymous):

im assuming by do you mean isolating y in the form of y=mx+b

OpenStudy (devan):

i mean solve

OpenStudy (devan):

for y

OpenStudy (amistre64):

ah, then abettu got it

OpenStudy (devan):

so would the answer be -1/2

OpenStudy (amistre64):

no, the answer is an equation for a line.

OpenStudy (devan):

but that's not and answer choice..